OpenStudy (anonymous):

if x and y are integers such that x^y=1, and xy=1, then y^x=? -1or 0 or 1 or 1.5 or 2

OpenStudy (mathstudent55):

Since xy = 1, neither x nor y can be zero.

OpenStudy (mathstudent55):

x^y = 1, but y is not zero, so x must be 1.

OpenStudy (mathstudent55):

Since x = 1, and xy = 1, y must be 1.

OpenStudy (mathstudent55):

Therefore y^x = 1.

OpenStudy (anonymous):

yes, x and y both are 1. I tried to solve it algebrically, using indices rules. thank u all

OpenStudy (anonymous):

mathstudent55, can u solve it algebrically?

OpenStudy (mathstudent55):

Algebraically: Since xy = 1, x and y are reciprocals. Therefore, y = 1/x Substituting in x^y = 1, we get \( \Large x^{\frac{1}{x}} = 1 \) \( \Large \log x^{\frac{1}{x}} = \log 1 \) \( \Large \dfrac{1}{x} \log x = 0 \) \(\large \log x = 0\) \(\large x = 1\) Once you have x = 1, y = 1/x = 1/1 = 1.

OpenStudy (anonymous):

thanks a lot.

OpenStudy (mathstudent55):

You're welcome.